# Self-generated electrokinetic flows from active-charged boundary patterns

### v1

@@ V Average unidirectional velocity
We find that the averaged velocity varies sinusoidally as function of the phase differences which is shown with the corresponding numerically obtained values in Fig. 3(d), for the case of equal ionic flux to diffusivity ratio $j_{0+}/D_{+}=j_{0-}/D_{-}$. The peak magnitude and direction reversing of the net velocity here are different depending phase difference combinations. For instance with $\alpha_{+}=0$, in the case when $\alpha_{-}=\pi/2$ the velocity goes as $\sim\sqrt{2}\sin(\beta+\pi/4)$ and in contrast for $\alpha_{-}=\pi$ it goes as $\sim 2\sin\beta$, where their magnitudes differ by a factor of $\sqrt{2}$ and direction switching happens at a distinct $\beta$ values (Fig. 3(d)). This means that in this case even when the relative cationic-anionic flux magnitudes are equal, there is an optimal pattern configuration that gives the maximal magnitude of unidirectional velocity. In our discussion earlier in Sec. IVB, we found that the cationic and anionic fluxes are need to be offset by some $\alpha_{-}\neq 2n\pi$ to give rise to non-vanishing unidirectional flow, and here we add that when they are particularly in the opposite phase $\alpha_{-}=(2n+1)\pi$ its magnitude is maximized. This configuration can be though of as a situation where the site of cationic in-flux and anionic out-flux coincide, or in another sense, there are biased active sites that allows for ion-selective fluxes. In this optimal case the effect coming from the cationic flux-charge phase difference are exactly superimposed with the aligning effect of the anionic counterpart, and thus the unidirectional flow is amplified.
We now estimate the magnitude of this net flow velocity in Eq. (33) for the case of only cationic flux active, by expressing its peak value as
$\langle v_{z}\rangle\sim g(j_{0+},\sigma_{0})\ R^{2}l\ \left(\frac{\lambda}{R}\right)^{3}\mathscr{L}_{\textrm{cyl}}(R/\lambda)$ (35)
@@
$g(j_{0+},\sigma_{0})=\frac{j_{0+}\sigma_{0}e}{2\pi\varepsilon\eta D_{+}}$ (36)
Table IV.2 shows the order of magnitude estimates for this speed at various ranges of parameter values. In particular, we use here the values for the amplitudes of cationic flux and surface charge $\{j_{0+},\sigma_{0}\}$ such that they are relatively small, meaning the relative (dimensionless) amplitudes $\delta_{+,\sigma}$ given by Eq. (18) are below 1. Under this criteria, in a micro-/nano-confinement we find when the Debye length is smaller (or bulk concentration is higher), larger values of flux and charge amplitudes can be adopted and as a consequence yields larger speeds. For example, at radius $R\sim 0.1$ $\mu$m and wavelength $l\sim 10$ $\mu$m, we get $R^{2}l\sim 0.1$ $\mu$m3 and when $\lambda\sim 10$ nm we have $\lambda/R\sim 0.1$ and $\mathscr{L}_{\textrm{cyl}}(R/\lambda)\sim 1$. If we use flux and charge amplitudes such that $\delta_{+,\sigma}\sim 10^{-3}$ with respect to this Debye length, we get $g\sim 10^{16}$ m-2 sec-1 leading to speeds like $\langle v_{z}\rangle\sim 1$ $\mu$m sec-1. At a smaller Debye length of $\lambda\sim 1$ nm (and thus now $\lambda/R\sim 10^{-2}$), if we keep the same (absolute) amplitudes of flux and charge the speed drops by three orders of magnitude. However, if instead use amplitudes that are relatively small now with respect to the smaller Deybe length then we get larger speeds like $\langle v_{z}\rangle\sim 10$ $\mu$m sec-1, and with much stronger flux it can be even larger going to $\sim 10$ cm sec-1. Thus, given such Debye lengths and amplitudes of ionic flux and surface charge, we can get here average (peak) flow speeds ranging from orders of microns per second up to tens of centimeters per second.
@@
Figure 4: Effect of ionic diffusivity mismatch. The stream plots of steady state fluid velocity in (r,z) obtained analytically from Eq. (29) and (30) with different ion diffusivity $D_{+}\neq D_{-}$ mismatch with $\alpha_{+}=0$ for (a) $(\alpha_{-},\beta)=(0,\pi/2)$ and (b) $(\alpha_{-},\beta)=(\pi/2,\pi/2)$. The corresponding modulation forms of the cationic (+) flux, the anionic (-) flux and the surface charge are shown above. The arrows of streamlines indicate the flow direction and the color map indicates the velocity magnitude given in terms of the characteristic flow speed $v_{0}$. (c) Plot of the wavelength averaged z-component velocity as a function of flux-charge phase difference $\beta$ at $(\alpha_{+},\alpha_{-})=(0,\pi/2)$, comparing the analytical approximation given by Eq. (33) with the numerical computed values (using COMSOL, Appendix C) and leading to an good agreement. (d) The landscape (contour plot) of normalized average velocity $\bar{\langle v_{z}\rangle}=\langle v_{z}\rangle/|\langle v_{z}\rangle|_{\text{max}}$ as function of $(\alpha_{-},\beta)$ at $\alpha_{+}=0$ for different diffusivity mismatch. The quantities are given in terms of the Debye length $\lambda$ for bulk concentration $C_{\infty}=1$ mM. We use here $R=10\lambda$ and $l=200\lambda$ for ions with same flux amplitudes $j_{0+}=j_{0-}$, and we use $D_{-}=5D_{+}$ for $D_{+}<D_{-}$ and $D_{+}=5D_{-}$ for $D_{+}>D_{-}$.
VI Effect of ion diffusivity mismatch
We examine in this section how changes in key physical parameters affects flow field and net velocity. We focus here on the effect of having mismatch in diffusivity of the ionic species, $D_{+}\neq D_{-}$ and compare the two cases $D_{+}<D_{-}$ and $D_{+}>D_{-}$. First we note that, in contrast to the case discussed in Sec. IVB, when a difference in ionic diffusivity (or more generally a difference in ionic flux-diffusivity ratio $j_{0+}/D_{+}\neq j_{0-}/D_{-}$) is introduced even when the cationic and anionic fluxes are in-phase, like at $\alpha_{-}=0$, a unidirectional flow emerges when offset by the flux-charge phase difference. Its magnitude and direction depends on the difference in the flux diffusivity ratio $(j_{0+}D_{-}-j_{0-}D_{+})/D_{+}D_{-}$ as can be seen from Eq. (33) for fluxes of same phase $\alpha_{-}=2n\pi$. If here we go from $D_{+}<D_{-}$ case to $D_{+}>D_{-}$ case, or vice versa, the unidirectional component must flip sign and hence the net flow reverses in direction like shown in Fig. 4(a). This means that the counteracting contributions from corresponding flux-charge phase differences are mismatch and subtracted with non-vanishing residue with a sign biasing the direction. However, for fluxes of opposite phase $\alpha_{-}=(2n+1)\pi$ they add together and the magnitude instead depends on the ratio $(j_{0+}D_{-}+j_{0-}D_{+})/D_{+}D_{-}$. Thus, switching the ionic diffusivity in this case may change the magnitude but does not affect the sign of the resulting unidirectional part and the net flow direction is preserved.
@@ VI Effect of ion diffusivity mismatch
In a physical sense, difference in diffusivity of the cationic and anionic species creates a lag in propagation that leads to a charge separation in the fluid. In regions away from the surface where the charge separation persists, additional electric fields are induced in response to mitigate this separation and bring it close to neutrality in the bulk. Near the surface, this can incur additional mismatched ionic gradients. When diffusivity is switched the propagation lag is flipped, reversing the charge separation and the direction of the induced fields are also reversed. Hence, the flow is affected by the overall contribution of these additional fields and gradients induced via propagation lag between the ionic species. In a mathematical sense, although our steady state Nernst-Planck equations in Eq. (6) are independent of the diffusivity coefficients, the ionic flux boundary conditions in Eq. (10) however are not. Hence, the diffusivity of the cationic and anionic species also controls the relative flux amplitudes $\delta_{\pm}$ of Eq. (18), which the coefficients $h_{0,c,s}$ in Eq. (21)-(23) that determine the leading flow velocity relies upon. Taken together, this effect of diffusivity mismatch can be seen as an attribute of a lag in ionic propagation that manifest in the steady state as a difference in the effective flux magnitudes which affects flow. This mismatch acts as an ion-specific parameter that controls the flow with the ability to induce velocity reversal as well as change of flow state.
VII Conclusion and outlook
@@
In this paper we established a surface-driven electrokinetic effect arising from active-charged patterns on confinement boundaries leading to directed electrolyte flow. Employing a continuum hydrodynamic description, we performed analytical approximations and validated with numerical finite element simulations. We show that modulation patterns of ionic fluxes and surface charge can self-generate a spectrum of flow states with a circulatory component, and notably, a unidirectional component that emerge when the flux and charge boundary patterns are offset by a flux-charge phase difference associated with the spatial mismatch between sites of peak flux and charge distributions. This can be achieved with just one ionic (cationic or anionic) flux active or with both fluxes active and exists even in the absence of externally applied fields or gradients in the bulk. We find selection rules with a rich dependence on the phase differences and amplitudes of fluxes and charge patterns lead the various flow states. When strong ionic flux constantly disrupts a tightly bound double layer, emergent lateral body forces become prominent driving fast flows. Varying key physical parameters such as the ionic diffusivity also affects flow fields and net velocity, resulting in velocity reversal and switching of the flow state. Taken together, our work provides a theoretical framework for a boundary-driven mechanism based on non-uniform patterning of ionic activity and surface charges, that self-generates directed flows in confined environments.
@@
Cellular boundaries often involve active and charged biological structures. To shed light on the underlying physical mechanisms behind various cellular processes, it is important to understand the role of boundary features that can regulate fluid motion in micro- and nano-confinements. At the same time, current developments in nanofluidic and iontronic technologies based on biological or soft matter systems, require control parameters that can functionalize the flow of confined electrolytes or ionic solutions. Our work here demonstrates a non-equilibrium steady state behavior of active-charged patterns inside capillaries that produce directed electrokinetic flows and gives quantitative insights on how such surface heterogeneties affect the state and magnitude of flow. Moreover, this work can be extended for systems of multiple active-charged channels or capillaries connected with each other, forming networks or circuits [15, 14]. For instance, in the circuit connections employed in nanofluidic or iontronic devices comprised of series, parallel or various combined configurations involving generalized network interactions of such capillary generators. Questions regarding the emerging physical principles governing active-charged networks can be further explored. In particular, those factors that determine fluid flow and ionic conductance for different network configurations, and tunable specifications that lead to optimal transport.
@@ Acknowledgements.
APPENDIX A LEADING-ORDER EQUATIONS
We give in this section the dimensionless forms of the reduced steady state governing equations with boundary conditions, and then write the leading order set of equations in the small-field and long wavelength approximations. The dimensionless transformations mentioned in Sec. III.2 turns Eqs. (4) and (6) into
$\nabla^{2}\tilde{\phi}=-\tilde{\rho}\,,$ (39)
$\nabla^{2}\tilde{\rho}+\nabla(\tilde{s}\nabla\tilde{\phi})=0\,,$ (40)
$\nabla^{2}\tilde{s}+\nabla(\tilde{\rho}\nabla\tilde{\phi})=0\,,$ (41)
@@
$(\nabla^{2}-1)\delta\tilde{\rho}=0\,,$ (45)
$\nabla^{2}\delta\tilde{\phi}=-\delta\tilde{\rho}\,,$ (46)
@@
and can be solved by separation of variable method [56]. This gives the analytical expressions given in Eqs (11) and (12) in the small-field approximations, that is when $\delta c_{\pm}\ll C_{\infty}/2$ or $\delta\tilde{\rho}\ll 1$, such that for $\tilde{R}\sim 1-10^{2}$ and $\tilde{k}\sim 10^{-5}-10^{-2}$,
$\frac{I_{0}(\sqrt{1+\tilde{k}^{2}}\tilde{R})\ \delta_{\pm,\sigma}}{\sqrt{1+\tilde{k}^{2}}\ I_{1}(\sqrt{1+\tilde{k}^{2}}\tilde{R})}\approx\delta_{\pm,\sigma}\ll 1$ (49)
Then to obtain the expression for the body force the respective $\tilde{r},\tilde{z}$ derivatives of Eqs (11) and (12) are taken. The resulting curl of this body force is now plugged into the right hand side of Eq. (25) giving
@@
Figure 5: Top panel: Schematic cross-section of the set-up employed in the numerical scheme. A cylindrical capillary of radius $R=0.1$ $\mu$m and length $L=18$ $\mu$m connecting two identical reservoirs make up the capillary-reservoir system of size $W=20$ $\mu$m by $H=1$ $\mu$m. Shown here are the boundary conditions applied at different regions of the walls enclosing the electrolyte. At the capillary surface the conditions are given by Eqs. (9) and (10), while connecting regions are set to satisfy the zero normal electric field and zero normal flux conditions. The far reservoir ends have conditions of zero electric potential and concentrations set to its bulk value. We impose no-slip conditions for the velocity on the walls. Bottom panel: Mesh grid pattern (zoomed in view of right hand section) implemented in COMSOL Multiphysics software [57] for numerically convergent solutions. The mesh sizes are adjusted (for $\lambda=0.1R$) with a finer mesh of $2\times 10^{-4}R$ near the wall and sparser mesh of $2\times 10^{-2}R$ away from it in the capillary domain, and then to larger sizes up to $0.1H$ in the reservoirs.
APPENDIX B STREAMFUNCTION COEFFICIENTS
@@
APPENDIX C NUMERICAL FINITE ELEMENT SIMULATIONS
@@
In this section we supplement the details about the scheme employed for the numerical solution of the steady state PNPNS system and the averaged velocity. We perform finite element simulations using the COMSOL Multiphysics software [57]. As in the main text, we use the time-independent PNPNS equations reduced at the low Péclet and low Reynolds number regime, leaving the set of Eqs. (4)-(8). This coupled system is solved in a cylindrical geometry for the set-up displayed in Fig. 5. Differences between the flow velocity obtained numerically and the analytically, are attributed to the small-field and long-wavelength approximations employed to obtain the analytical solutions and the finite length set-up implemented for the numerical simulation. However, despite these differences the results agree well as shown in Fig. 3 of the main text.

### v2

@@ V Average unidirectional velocity
We find that the averaged velocity varies sinusoidally as function of the phase differences which is shown with the corresponding numerically obtained values in Fig. 3(d), for the case of equal ionic flux to diffusivity ratio $j_{0+}/D_{+}=j_{0-}/D_{-}$. The peak magnitude and direction reversing of the net velocity here are different depending on phase difference combinations. For instance with $\alpha_{+}=0$, in the case when $\alpha_{-}=\pi/2$ the velocity goes as $\sim\sqrt{2}\sin(\beta+\pi/4)$ and in contrast for $\alpha_{-}=\pi$ it goes as $\sim 2\sin\beta$, where their magnitudes differ by a factor of $\sqrt{2}$ and direction switching happens at a distinct $\beta$ values (Fig. 3(d)). This means that in this case even when the relative cationic-anionic flux magnitudes are equal, there is an optimal pattern configuration that gives the maximal magnitude of unidirectional velocity. In our discussion earlier in Sec. IVB, we found that the cationic and anionic fluxes are need to be offset by some $\alpha_{-}\neq 2n\pi$ to give rise to non-vanishing unidirectional flow, and here we add that when they are particularly in the opposite phase $\alpha_{-}=(2n+1)\pi$ its magnitude is maximized. This configuration can be thought of as a situation where the site of cationic in-flux and anionic out-flux coincide, or in another sense, there are biased active sites that allows for ion-selective fluxes. In this optimal case the effect coming from the cationic flux-charge phase difference is exactly superimposed with the aligning effect of the anionic counterpart, and thus the unidirectional flow is amplified.
We now estimate the magnitude of this net flow velocity based on our analytic approximation for the case of only cationic flux active ($j_{0-}=0$) in Eq. (33), by expressing its peak value ($\beta-\alpha_{+}=\pi/2$) as
$\langle v_{z}\rangle\approx\delta_{+}\delta_{\sigma}\ V$ (35)
@@
$V=\frac{k_{B}TC_{\infty}\lambda}{\eta}\ \frac{l}{2\pi R}\ \mathscr{L}_{\textrm{cyl}}(R/\lambda)$ (36)
In particular, we use the values for the amplitudes of cationic flux and surface charge such that they are relatively small in the long-wavelength limit, meaning the relative (dimensionless) amplitudes $\delta_{+,\sigma}<1$ and $kR>2\delta_{+}$ (see Appendix B). Under these criteria, when the Debye length is smaller (or bulk concentration is higher), larger values of flux and charge amplitudes can be adopted yielding larger speeds. For example, at radius $R=0.5$ $\mu$m and wavelength $l=1$ $\mu$m, when $\lambda\approx 14$ nm we have $\mathscr{L}_{\textrm{cyl}}(R/\lambda)\approx 0.85$. We get the limiting speed $V\approx 0.1$ m s-1 leading to the average unidirectional speeds like $\langle v_{z}\rangle\approx 1$ mm s-1. At even smaller Debye lengths, if we keep the same (absolute) amplitudes of flux and charge the speed drops by three orders of magnitude. However, if instead use amplitudes that are relatively small now with respect to the smaller Debye length then we get larger speeds, and with much stronger flux. Table IV.2 shows the estimates for this speed at various ranges of parameter values. Thus, given such Debye lengths and amplitudes of ionic flux and surface charge, we can get here average (peak) flow speeds ranging from orders of microns up to millimeters per second.
Figure 4: Effect of ionic diffusivity mismatch. The stream plots of steady state fluid velocity in $(\tilde{r},\tilde{z})$ obtained analytically from Eq. (29) and (30) with different ion diffusivity $D_{+}\neq D_{-}$ mismatch with $\alpha_{+}=0$ for (a) $(\alpha_{-},\beta)=(0,\pi/2)$ and (b) $(\alpha_{-},\beta)=(\pi/2,\pi/2)$. The corresponding modulation forms of the cationic (+) flux, the anionic (-) flux and the surface charge are shown above. The arrows of streamlines indicate the flow direction and the color map indicates the velocity magnitude given in dimensionless units as in Fig. 2. (c) Plot of the wavelength averaged z-component velocity as a function of flux-charge phase difference $\beta$ at $(\alpha_{+},\alpha_{-})=(0,\pi/2)$ where $\langle\bar{v}_{z}\rangle=\langle v_{z}\rangle/\lambda$ s-1 as in Fig. 3, comparing the analytical approximation given by Eq. (33) with the numerical computed values (using COMSOL, Appendix D) and leading to an good agreement. (d) The landscape (contour plot) of normalized average velocity $\bar{\langle v_{z}\rangle}=\langle v_{z}\rangle/|\langle v_{z}\rangle|_{\text{max}}$ as function of $(\alpha_{-},\beta)$ at $\alpha_{+}=0$ for different diffusivity mismatch. We use here $R=10\lambda$ and $l=200\lambda$ for ions with same flux amplitudes $j_{0+}=j_{0-}$, and we use $D_{-}=5D_{+}$ for $D_{+}<D_{-}$ and $D_{+}=5D_{-}$ for $D_{+}>D_{-}$.
@@
VI Ion diffusivity mismatch
We examine in this section how changes in key physical parameters affects flow field and net velocity. We focus here on the effect of having mismatch in diffusivity of the ionic species, $D_{+}\neq D_{-}$ and compare the two cases $D_{+}<D_{-}$ and $D_{+}>D_{-}$. First we note that, in contrast to the case discussed in Sec. IVB, when a difference in ionic diffusivity (or more generally a difference in ionic flux-diffusivity ratio $j_{0+}/D_{+}\neq j_{0-}/D_{-}$) is introduced even when the cationic and anionic fluxes are in-phase, like at $\alpha_{-}=0$, a unidirectional flow emerges when offset by the flux-charge phase difference.
Along with the magnitude of the net flow velocity, its direction also depends on the difference in the diffusivity ratio $(j_{0+}D_{-}-j_{0-}D_{+})/D_{+}D_{-}$ as can be seen from Eq. (33) for fluxes of same phase $\alpha_{-}=2n\pi$. If we switch here from the $D_{+}<D_{-}$ case to the $D_{+}>D_{-}$ case, or vice versa, the unidirectional component must flip sign and hence the net flow reverses in direction like shown in Fig. 4(a). This means that the counteracting contributions from corresponding flux-charge phase differences are mismatch and subtracted with non-vanishing residue with a sign biasing the direction. However, for fluxes of opposite phase $\alpha_{-}=(2n+1)\pi$ they add together and the magnitude instead depends on the ratio $(j_{0+}D_{-}+j_{0-}D_{+})/D_{+}D_{-}$. Thus, switching the ionic diffusivity in this case may change the magnitude but does not affect the sign of the resulting unidirectional part and the net flow direction is preserved.
Figure 5: Enzyme powered flows generated by flux-charge patch separation and enhancement from diffusivity mismatch in microfluidic chamber. (a) Illustration of the zoomed-in view as in Fig. 1 of the interface between the capillary walls and the liquid solution inside showing a flux patch (green) and a charge patch (purple) separated by distance $\Delta$. The stream plots of steady state fluid velocity in (r,z) obtained numerically (using COMSOL, Appendix D) for (b) $\Delta=0$ and (c) $\Delta=1$ $\mu$m with different ion diffusivity mismatch. The corresponding Gaussian forms of the cationic (+) flux, the anionic (-) flux and the surface charge are shown above. The arrows of streamlines indicate the flow direction and the color map indicates the velocity magnitude. (d) Plot of average unidirectional velocity as a function of patch separation for different diffusivity mismatch. We use here bulk concentration $C_{\infty}=0.1$ mM, capillary radius $R=0.5$ $\mu$m, $j_{0}=5\times 10^{-4}$ mM m s-1, $\sigma_{0}=-1\times 10^{-5}$ C m-2, and Gaussian spread factor $\xi=2$ $\mu$m-2.
@@ VI Ion diffusivity mismatch
In a mathematical sense, although our steady state Nernst-Planck equations in Eq. (6) are independent of the diffusivity coefficients, the ionic flux boundary conditions in Eq. (10) however are not. Hence, the diffusivity of the cationic and anionic species also controls the relative flux amplitudes $\delta_{\pm}$ of Eq. (18), which the coefficients $h_{0,c,s}$ in Eq. (21)-(23) that determine the leading flow velocity rely upon. In a physical sense, difference in diffusivity of the cationic and anionic species creates a charge separation in the fluid at the steady state. In regions away from the surface where the charge separation persists, additional electric fields are induced in response to mitigate this separation and bring it close to neutrality in the bulk. Near the surface, this can incur additional mismatched ionic gradients. When diffusivity is switched, the charge separation reverses and the direction of the induced fields are hence flipped. The flow is affected by the overall contribution of these additional fields and gradients induced separation between the ionic species. Taken together, this effect of diffusivity mismatch can be an attribute of the additional gradients induced at the steady state as a consequence of difference in the effective flux magnitudes, which in turn affects the flow. This mismatch acts as an ion-specific parameter that controls the flow with the ability to optimize the speed as well as induce velocity reversal or change of flow state.
VII Enzyme powered microfluidic generator
[…]
VII.1 Gaussian patch model
We consider a finite sized axisymmetric system where a cylindrical capillary of diameter $1$ $\mu$m ($R=0.5$ $\mu$m) and length $10$ $\mu$m connects two large electrolyte reservoirs (see Appendix D). At the steady state, we assume wall patch distributions of ionic flux and charge density takes a Gaussian form given as follows
$j_{\pm}(z)=j_{0}e^{-\xi(z-\Delta/2)^{2}}\quad\textrm{at}\quad r=R$ (39)
$\sigma(z)=\sigma_{0}e^{-\xi(z+\Delta/2)^{2}}\quad\textrm{at}\quad r=R$ (40)
where the Gaussian spread factor $\xi$ characterizes the patch size and the flux-charge separation distance $\Delta$ characterizes the displacement between sites of the peak densities. Based on the previous work [37], the magnitude of cationic and anionic flux settles to $j_{0}\sim 10^{-4}$ mM m s-1 (in-flux for urease) and the surface charge to $\sigma_{0}\sim 10^{-5}$ C m-2 (negative for silica). These are micron sized patches corresponding to $\xi\sim 1$ $\mu$m-2 and we vary $\Delta$ from 00 up to $\sim 5$ $\mu$m (see top panels of Fig. 5(b) and (c)).
VII.2 Flows generated by flux-charge patch
In panels (b)-(d) of Fig. 5, we display the numerical results obtained for this patch model. When $\Delta=0$ there is complete overlap between the patch of flux and charge and leads to a circulatory flow at the site of peak density (Fig. 5(b)). However, when they are displaced with a nonzero separation between their peak sites such as $\Delta=1$ $\mu$m, a unidirectional component develops and gives rise to localized vortices near at the site of peak charge (Fig. 5(c)). These spots have velocity profiles representative of the intermediate state as shown in Fig. 2(a) of $U\pm W$ corresponding to the different cases of diffusivity mismatch. We find that the magnitude of the directed flow increase with larger separation distance and is enhanced for the case with larger diffusivity mismatch (Fig. 5(d)). We note that this net flow velocity can be further optimized by grafting a patterned sequence of multiple pairs of such flux-charge patches.
VIII Conclusions and outlook
@@
In this paper we established a boundary-driven electrokinetic effect arising from active-charged patterns on confinement walls leading to directed electrolyte flow. Employing a continuum hydrodynamic description, we performed analytical approximations and validated with numerical finite element simulations. We show that modulation patterns of ionic fluxes and surface charge can self-generate a spectrum of flow states with a circulatory component, and notably, a unidirectional component that emerge when the flux and charge boundary patterns are offset by a flux-charge phase difference associated with the spatial mismatch between sites of peak flux and charge distributions. This can be achieved with just one ionic (cationic or anionic) flux active or with both fluxes active and exists even in the absence of externally applied fields or gradients in the bulk. We find selection rules with a rich dependence on the phase differences and amplitudes of fluxes and charge patterns lead the various flow states. When strong ionic flux constantly disrupts a non-uniform double layer, emergent lateral body forces become prominent driving fast flows. Varying key physical parameters such as the ionic diffusivity also affects flow fields and net velocity, resulting in velocity reversal, switching of the flow state and speed enhancement. Taken together, our work provides a theoretical framework for a boundary-driven mechanism based on non-uniform patterning of ionic activity and surface charges, that self-generates directed flows in confined environments.
@@
Cellular boundaries often involve active and charged biological structures. To shed light on the underlying physical mechanisms behind various cellular processes, it is important to understand the role of boundary features that can regulate fluid motion in micro- and nano-confinements [62, 63, 64]. At the same time, current developments in fluidic and iontronic technologies based on biological or soft matter systems, require control parameters that can functionalize the flow of confined electrolytes or ionic solutions. Our work here demonstrates a non-equilibrium steady state behavior of active-charged patterns inside capillaries that produce directed electrokinetic flows and gives quantitative insights on how such surface heterogeneties affect the state and magnitude of flow. Moreover, this work can be extended for systems of multiple active-charged channels or capillaries connected with each other, forming networks or circuits [15, 14]. For instance, in the circuit connections employed in fluidic or iontronic devices comprised of series, parallel or various combined configurations involving generalized network interactions of such capillary generators. Questions regarding the emerging physical principles governing active-charged networks can be further explored. In particular, those factors that determine fluid flow and ionic conductance for different network configurations, and tunable specifications that lead to optimal transport.
@@ Acknowledgements.
APPENDIX A LEADING ORDER EQUATIONS
We give in this section the dimensionless forms of the reduced steady state governing equations with boundary conditions, and then write the leading order set of equations in the small-amplitude and long-wavelength approximations. The dimensionless transformations mentioned in Sec. III.2 turns Eqs. (4) and (6) into
~$\tilde{\nabla}^{2}\tilde{\phi}=-\tilde{\rho}\,,$ (41)
~$\tilde{\nabla}^{2}\tilde{\rho}+\nabla(\tilde{s}\tilde{\nabla}\tilde{\phi})=0\,,$ (42)
~$\tilde{\nabla}^{2}\tilde{s}+\tilde{\nabla}(\tilde{\rho}\tilde{\nabla}\tilde{\phi})=0\,,$ (43)
@@
(~$(\tilde{\nabla}^{2}-1)\delta\tilde{\rho}=0\,,$ (47)
~$\tilde{\nabla}^{2}\delta\tilde{\phi}=-\delta\tilde{\rho}\,,$ (48)
@@
and can be solved by separation of variable method [65]. This gives the analytical expressions given in Eqs (11) and (12) in the small-amplitude approximations. Then to obtain the expression for the body force the respective $\tilde{r},\tilde{z}$ derivatives of Eqs (11) and (12) are taken. The resulting curl of this body force is now plugged into the right hand side of Eq. (25) giving
@@
Figure 6: Log-log plot of average unidirectional velocity as a function of relative flux-charge amplitudes $\delta_{+,\sigma}$ at different $kR$. The analytical expression Eq. (57) is compared with the numerical results from COMSOL, and $V$ is given by Eq. (36). We use here $C_{\infty}=1$ mM and $R=0.5$ $\mu$m.
APPENDIX B SMALL-AMPLITUDE AND LONG-WAVELENGTH APPROXIMATION
In this section, we look at the limits for which the small-amplitude and long-wavelength approximations hold. The ionic fluxes and surface charge are taken to be small such that the perturbed concentration fields are $\delta c_{\pm}\ll C_{\infty}/2$ or $\delta\tilde{\rho}\ll 1$ which gives in the long wavelength limit,
$\frac{I_{0}(\sqrt{1+\tilde{k}^{2}}\tilde{R})}{\sqrt{1+\tilde{k}^{2}}\ I_{1}(\sqrt{1+\tilde{k}^{2}}\tilde{R})}\delta_{\pm,\sigma}\approx\delta_{\pm,\sigma}\ll 1$ (55)
Along with this, the magnitude of the electric field is also considered small, that is |~$|\tilde{\nabla}\delta\tilde{\phi}|\ll 1$. The $r$-component ($\partial_{\tilde{r}}\tilde{\phi}$) gives a similar condition $\delta_{\pm,\sigma}\ll 1$ and the $z$-component ($\partial_{\tilde{z}}\tilde{\phi}$) we have again $\delta_{\sigma}\ll 1$ but also
$\left|\frac{I_{0}(\tilde{k}\tilde{R})}{I_{1}(\tilde{k}\tilde{R})}-\frac{\tilde{k}\ I_{0}(\sqrt{1+\tilde{k}^{2}}\tilde{R})}{\sqrt{1+\tilde{k}^{2}}\ I_{1}(\sqrt{1+\tilde{k}^{2}}\tilde{R})}\right|\ \delta_{\pm}\approx\frac{2\delta_{\pm,\sigma}}{\tilde{k}\tilde{R}}\ll 1$ (56)
This implies restrictions on wavelength and radius such that $\tilde{k}\tilde{R}>2\delta_{\pm}$. For example, if $l\approx 20R$ so that $kR=2\pi R/l\approx 0.3$ then $\delta_{\pm}\lesssim 0.15$, or conversely if $\delta_{\pm}\approx 0.1$, then $l\lesssim 30R$.
We compare the analytical and numerical average velocity for the case $\delta_{+}\approx\delta_{\sigma}$ where the analytical expression of Eq. (35) written in log-log form is
$\log_{10}[\langle v_{z}\rangle/V]\approx 2\log_{10}[\delta_{+,\sigma}]$ (57)
where $V$ is given by Eq. (36). Fig. 6 shows the log-log plot comparing the analytical and numerical values, where we find an excellent matching for $\delta_{+,\sigma}\lesssim 0.1$ with $kR\gtrsim 2$.
Figure 7: Top panel: Schematic cross-section of the set-up employed in the numerical scheme. A cylindrical capillary of radius $R$ and length $L$ connecting two identical reservoirs make up the capillary-reservoir system of size $W$ by $H$. Shown here are the boundary conditions applied at different regions of the walls enclosing the electrolyte. At the capillary surface the conditions are given by Eqs. (9) and (10), while connecting regions are set to satisfy the zero normal electric field and zero normal flux conditions. We impose no-slip and no-penetration conditions for the velocity on the walls. The far reservoir ends have conditions of zero electric potential and concentrations set to its bulk value and zero velocity. Bottom panel: Mesh grid pattern (zoomed in view of right hand section) implemented in COMSOL Multiphysics software [66] for numerically convergent solutions. The mesh sizes are adjusted (for $\lambda=0.1R$) with a finer mesh of $2\times 10^{-4}R$ near the wall and sparser mesh of $2\times 10^{-2}R$ away from it in the capillary domain, and then to larger sizes up to $0.1H$ in the reservoirs.
APPENDIX C STREAMFUNCTION COEFFICIENTS
@@
APPENDIX D NUMERICAL FINITE ELEMENT SIMULATIONS
@@
In this section we supplement the details about the scheme employed for the numerical solution of the steady state PNPNS system and the averaged velocity. We perform finite element simulations using the COMSOL Multiphysics software [66]. As in the main text, we use the time-independent PNPNS equations reduced at the low Péclet and low Reynolds number regime, leaving the set of Eqs. (4)-(8). This coupled system is solved in a cylindrical geometry for the set-up displayed in Fig. 7. Differences between the flow velocity obtained numerically and the analytically, are attributed to the small-amplitude and long-wavelength approximations employed to obtain the analytical solutions and the finite length set-up implemented for the numerical simulation. However, despite these differences the results agree well as shown in Fig. 3 of the main text and in Fig. 5 of Appendix B.